# Multi-intention Inverse Q-learning for Interpretable Behavior Representation

### v3

@@ C.1 Gridworld benchmark
All evaluated algorithms were fit to multiple datasets with different number of expert demonstrations, and each dataset analyzed using a $5$-fold cross-validation. Note that since the EM process does not guarantee to find the global optimum, for each training, HIAVI was randomly initialized repeatedly for $10$ times and the best performed initialization (evaluated by test set log-likelihood) was selected for analysis. The initial intention distribution $\Pi$ was initialized uniformly, and the intention transition matrix $\Lambda$ was initialized as: $\Lambda=0.95\times I+\mathcal{N}(0,0.05\times I)$, where $\mathcal{N}$ denotes the normal distribution and $I\in\mathbf{R}^{2\times 2}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to $1$. The discount factor of the MDP was set to be $\gamma=0.9$.
@@ C.2 Real-world mice navigation benchmark
For comparability with the results in [AJP22], we obtained their pre-processed mouse trajectories for water-restricted and water-unrestricted animals from https://github.com/97aditi/dynamic_irl. The original recorded animal trajectories from [RZPM21] are provided with MIT open source license at the following repository: https://github.com/markusmeister/Rosenberg-2021-Repository. For the pre-processing, Ashwood et al. [AJP22] used a clustering algorithm (based on DBSCAN [EKS+96]) for aligning trajectories across animals and bouts to reduce variability. After the pre-processing, they obtained $200$ trajectories from the water-restricted animals and $207$ trajectories from the water-unrestricted animals. $20\%$ of trajectories from each cohort were held out as a test set.
We used the source code and the best performed set of hyperparameters provided by [AJP22] to train DIRL on the animal trajectories. All HIAVI algorithms were trained for $10$ repeated runs with different initializations, and the results from the initializations with the hightest test set log-likelihood was selected for analysis. The initial intention distribution $\Pi$ was initialized uniformly, and the intention transition matrix $\Lambda$ was initialized as: $\Lambda=0.95\times I+\mathcal{N}(0,0.05\times I)$, where $\mathcal{N}$ denotes the normal distribution and $I\in\mathbf{R}^{K\times K}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to $1$. The MDP was defined to have a deterministic state transition function $P$ and the discount factor was set to be $\gamma=0.99$.

### v4

@@ C.1 Gridworld benchmark
The whole expert demonstration dataset consisted of $1024$ trajectories. All evaluated algorithms were fit to multiple sub-datasets with different number of expert trajectories, and each dataset was analyzed using a $5$-fold cross-validation. Note that since the EM process does not guarantee to find the global optimum, for each training, HIAVI was randomly initialized repeatedly for $10$ times and the best performed initialization (evaluated by test set log-likelihood) was selected for analysis. The initial intention distribution $\Pi$ was initialized uniformly, and the intention transition matrix $\Lambda$ was initialized as: $\Lambda=0.95\times I+\mathcal{N}(0,0.05\times I)$, where $\mathcal{N}$ denotes the normal distribution and $I\in\mathbf{R}^{2\times 2}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to $1$. The discount factor of the MDP was set to be $\gamma=0.9$.
@@ C.2 Real-world mice navigation benchmark
For comparability with the results in Ashwood et al. (2022a), we obtained their pre-processed mouse trajectories for water-restricted and water-unrestricted animals from https://github.com/97aditi/dynamic_irl. The original recorded animal trajectories from Rosenberg et al. (2021) are provided with MIT open source license at the following repository: https://github.com/markusmeister/Rosenberg-2021-Repository. For the pre-processing, Ashwood et al. (2022a) used a clustering algorithm (based on DBSCAN (Ester et al., 1996)) for aligning trajectories across animals and bouts to reduce variability. After the pre-processing, they obtained $200$ trajectories from the water-restricted animals and $207$ trajectories from the water-unrestricted animals. $20\%$ of trajectories from each cohort were held out as a test set.
We used the source code and the best performed set of hyperparameters provided by Ashwood et al. (2022a) to train DIRL on the animal trajectories. All HIAVI algorithms were trained for $10$ repeated runs with different initializations, and the results from the initializations with the highest test set log-likelihood was selected for analysis. The initial intention distribution $\Pi$ was initialized uniformly, and the intention transition matrix $\Lambda$ was initialized as: $\Lambda=0.95\times I+\mathcal{N}(0,0.05\times I)$, where $\mathcal{N}$ denotes the normal distribution and $I\in\mathbf{R}^{K\times K}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to $1$. The MDP was defined to have a deterministic state transition function $P$ and the discount factor was set to be $\gamma=0.99$.